OpenStudy (anonymous):

Which ordered pair is a solution of the inequality? y4x - 5 A. (3, 4) B. (2, 1) C. (3, 0) D. (1, 1)

OpenStudy (anonymous):

y>4x-5

OpenStudy (anonymous):

hello again glad to see u

OpenStudy (cwrw238):

all you have to do here is plug in the 4 values of x and y and see if they fit

OpenStudy (anonymous):

im sorry idk what that means

OpenStudy (cwrw238):

eg A: x=3, y=4 substituting 3 > 4(3) -5 = 7 3 not > 7 so its not A

OpenStudy (anonymous):

i know its not b because i got b wrong on the test

OpenStudy (anonymous):

i think its c

OpenStudy (cwrw238):

try D instead of x put 1 and instead of y put 1 (in the formula y> 4x - 5

OpenStudy (anonymous):

3 > 4(1) -5 = 5

OpenStudy (anonymous):

did i do that right?

OpenStudy (cwrw238):

no 1 > 4(1) - 5 = 4 -5 = -1 so 1 > -1 is that right?

OpenStudy (anonymous):

yes?!

OpenStudy (cwrw238):

yes 1 is greater than -1 so the correct choice is D

OpenStudy (anonymous):

ooooh ok im staRTING TO GET IT LITTLE BY LITTLE i didnt get it back then in highschool thank god for smart people like u and this website when i came acrossed it i thought it was the best thing ever

OpenStudy (cwrw238):

no probs - glad to be of help